OpenStudy (anonymous):

Let \(a_{n}\) be the nth term of an arithmetic sequence. If \(a_{18}\)=26 and \(a_{23}\)=61, which of the following are true? A. \(a_{14}\)<0 B. \(a_{1}-a_{2}\)<0 C. \(a_{1}+a_{2}+a_{3}+...+a_{27}\)>0

OpenStudy (shubhamsrg):

you can surely find a and d from the given info ?

OpenStudy (anonymous):

how @@

OpenStudy (anonymous):

This type of question i don't know how to do .

OpenStudy (shubhamsrg):

You know the nth term of an arithmetic progression /arithmetic sequence/ A.P.

OpenStudy (shubhamsrg):

?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

a+19d=26 a+24d=61 a=-107 d=7

OpenStudy (shubhamsrg):

nop its a+ (n-1)d so your eqns will be a+17d = 26 and a+ 22d = 61

OpenStudy (anonymous):

OMG, messy formula..

OpenStudy (anonymous):

a=-93 d=7

OpenStudy (shubhamsrg):

right now it becomes easier to check which ones are true and false ?

OpenStudy (anonymous):

let me try

OpenStudy (anonymous):

\(a_{14}\)=-2

OpenStudy (shubhamsrg):

right

OpenStudy (anonymous):

\(a_{2}\)=-86 ?

OpenStudy (shubhamsrg):

correct

OpenStudy (anonymous):

\(a_{27}\)=89 \[S(27)=(\frac{ (-93+89)\times27 }{ 2 })=-54 ???\]

OpenStudy (anonymous):

A and B are correct then ?

OpenStudy (shubhamsrg):

yes, bingo!

OpenStudy (anonymous):

thank you very much

OpenStudy (shubhamsrg):

glad to help